Angle A and angle B are supplementary angles and angle A is 5° less than four times angle B. Find the measures of angle A and angle B
Answer by ikleyn(52807)

.

Your equation is


    B + (4B - 5) = 180  degrees.


    5B = 185

     B = 185/5 = 37°.


     A = 180 - 37 = 143°.


ANSWER.  A = 143°;  B = 37°.

Solved.

-----------------

See the lessons
    - Angles basics
    - Solved problems on supplementary and complementary angles
in this site.

They are written specially for you.

After reading these lessons you will tackle such problems very easy on your own, without asking for the help from outside.

Consider them as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.